# Gravitational radiation reaction around a static black hole surrounded by a Dehnen type dark matter halo

### v1

@@ III.2 Orbital parameters: $p$ and $e$
Using Eqs. (25)-(26) and (32)-(41), the rate of change of the eccentricity and semi-latus rectum are given by Cao et al. (2025)
$\frac{de}{dt}$ $=$ $\frac{e^{2}-1}{2e}\left(\frac{dE/dt}{E}+2\frac{dL/dt}{L}\right),$ (44)
$\frac{dp}{dt}$ $=$ $\frac{da}{dt}\left(1-e^{2}\right)-2ae\frac{de}{dt},$ (45)
@@ IV Gravitational waveforms
$h_{ij}=\frac{4G\eta\mathcal{M}}{c^{4}D_{L}}\left(v_{i}v_{j}-\frac{G\mathcal{M}}{r}n_{i}n_{j}\right),$ (46)
@@
$\mathbf{e}_{X}$ $=(\cos\zeta,-\sin\zeta,0),$ (47)
@@
$h_{+}$ $=-\frac{2\eta(G\mathcal{M})^{2}}{c^{4}D_{L}r}(1+\cos^{2}\iota)\cos(2\phi+2\zeta),$ (48)
@@
To assess the effect of a Dehnen DM halo on the evolution of EMRIs, we compute the number of orbital cycles accumulated during the inspiral as Maselli et al. (2020); Barsanti et al. (2022)
$\mathcal{N}=\frac{1}{\pi}\int_{t_{1}}^{t_{2}}\dot{\phi}(t)\,dt.$ (49)
Using Eqs. (27) - (28), and (49), we evaluate $\mathcal{N}(t)$ numerically for the EMRI model under consideration. Fig. 6 shows the resulting $\mathcal{N}(t)$ for EMRIs with and without the DM halo, indicating that the presence of DM produces a slight reduction in $\mathcal{N}(t)$ relative to the vacuum case. This effect becomes more pronounced for halos that are more concentrated (smaller $r_{s}$) and denser (larger $\rho_{s}$), as the additional gravitational potential and dynamical effects from the halo slow down the system evolution. These findings suggest that dense or extended DM environments around EMRIs may alter their GW signature.
V Final remarks
In this work, we have examined how the DM halo of the Dehnen type affects the dynamics and GW signatures of EMRIs. We have found that DM introduces systematic deviations from the vacuum case in all aspects of EMRI evolution which may be astrophysically important on long spiraling scales. We started with orbital period and pericenter precession, and we showed that both of these quantities are sensitive to the properties of the halo. Higher central density $\rho_{s}$ and greater radius of $r_{s}$ result in shorter orbital periods but more intense pericenteral precession $\Delta\phi$. Although variations in halo density $\rho_{s}$ cause slight changes, the scale radius has a stronger effect on orbital period at large distances and precession at small distances.
@@ V Final remarks
Acknowledgements.
A.A. thanks the University of Bologna theory group for their hospitality, and is particularly grateful to Silvia Pascoli for her support during his visit. The visit was funded by the European Union’s Horizon Europe research and innovation programme under Marie Skłodowska -Curie Staff Exchange grant agreement No 101086085-ASYMMETRY. He also acknowledges support from Mahmoud Safari and the Iranian Science Foundation, grant 4031449. R.C. is partially supported by the INFN grant FLAG, and his work has also been carried out in the framework of activities of the National Group of Mathematical Physics (GNFM, INdAM) and the COST Action CA23115 (RQI). O.L. acknowledges financial support from the Fondazione ICSC, Spoke 3 Astrophysics and Cosmos Observations. National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) Project ID CN_\_00000013 ”Italian Research Center on High-Performance Computing, Big Data and Quantum Computing” funded by MUR Missione 4 Componente 2 Investimento 1.4: Potenziamento strutture di ricerca e creazione di ”campioni nazionali di R&\&S (M4C2-19 )” - Next Generation EU (NGEU) GRAB-IT Project, PNRR Cascade Funding Call, Spoke 3, INAF Italian National Institute for Astrophysics, Project code CN00000013, Project Code (CUP): C53C22000350006, cost center STI442016.

### v2

@@ 3.2 Orbital parameters: $p$ and $e$
Using Eqs. (3.19)-(3.20) and (3.26)-(3.35), the rate of change of the eccentricity and semi-latus rectum are given by [14]
$\frac{de}{dt}$ $=$ $\frac{e^{2}-1}{2e}\left(\frac{dE/dt}{E}+2\frac{dL/dt}{L}\right),$ (3.38)
$\frac{dp}{dt}$ $=$ $\frac{da}{dt}\left(1-e^{2}\right)-2ae\frac{de}{dt},$ (3.39)
@@ 4 Gravitational waveforms
$h_{ij}=\frac{4G\eta\mathcal{M}}{c^{4}D_{L}}\left(v_{i}v_{j}-\frac{G\mathcal{M}}{r}n_{i}n_{j}\right),$ (4.1)
@@
$\mathbf{e}_{X}$ $=(\cos\zeta,-\sin\zeta,0),$ (4.2)
@@
$h_{+}$ $=-\frac{2\eta(G\mathcal{M})^{2}}{c^{4}D_{L}r}(1+\cos^{2}\iota)\cos(2\phi+2\zeta),$ (4.3)
@@
To assess the effect of a Dehnen DM halo on the evolution of EMRIs, we compute the number of orbital cycles accumulated during the inspiral as [38, 8]
$\mathcal{N}=\frac{1}{\pi}\int_{t_{1}}^{t_{2}}\dot{\phi}(t)\,dt.$ (4.4)
Using Eqs. (3.21) - (3.22), and (4.4), we evaluate $\mathcal{N}(t)$ numerically for the EMRI model under consideration. Fig. 6 shows the resulting $\mathcal{N}(t)$ for EMRIs with and without the DM halo, indicating that the presence of DM produces a slight reduction in $\mathcal{N}(t)$ relative to the vacuum case. This effect becomes more pronounced for halos that are more concentrated (smaller $r_{s}$) and denser (larger $\rho_{s}$), as the additional gravitational potential and dynamical effects from the halo slow down the system evolution. These findings suggest that dense or extended DM environments around EMRIs may alter their GW signature.
5 Final remarks
In this work, we have examined how the DM halo of the Dehnen type affects the dynamics and GW signatures of EMRIs. After establishing the theoretical motivation for adopting the Dehnen profile as the dark-matter background and outlining its key physical properties, we find that the presence of dark matter induces systematic deviations from the vacuum case across all aspects of EMRI evolution, potentially yielding astrophysically significant effects over long inspiral timescales. We started with orbital period and pericenter precession, and we showed that both of these quantities are sensitive to the properties of the halo. Higher central density $\rho_{s}$ and greater radius of $r_{s}$ result in shorter orbital periods but more intense pericenteral precession $\Delta\phi$. Although variations in halo density $\rho_{s}$ cause slight changes, the scale radius has a stronger effect on orbital period at large distances and precession at small distances.
@@ 5 Final remarks
Acknowledgments
A.A. thanks the University of Bologna theory group for their hospitality, and is particularly grateful to Silvia Pascoli for her support during his visit. The visit was funded by the European Union’s Horizon Europe research and innovation programme under Marie Skłodowska -Curie Staff Exchange grant agreement No 101086085-ASYMMETRY. He also acknowledges support from Mahmoud Safari and the Iranian Science Foundation, grant 4031449. R.C. is partially supported by the INFN grant FLAG, and his work has also been carried out in the framework of activities of the National Group of Mathematical Physics (GNFM, INdAM) and the COST Action CA23115 (RQI). O.L. acknowledges financial support from the Fondazione ICSC, Spoke 3 Astrophysics and Cosmos Observations. National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) Project ID CN_\_00000013 "Italian Research Center on High-Performance Computing, Big Data and Quantum Computing" funded by MUR Missione 4 Componente 2 Investimento 1.4: Potenziamento strutture di ricerca e creazione di "campioni nazionali di R&\&S (M4C2-19 )" - Next Generation EU (NGEU) GRAB-IT Project, PNRR Cascade Funding Call, Spoke 3, INAF Italian National Institute for Astrophysics, Project code CN00000013, Project Code (CUP): C53C22000350006, cost center STI442016.